# Distributed Scalable Coupled Policy Algorithm for Networked Multi-agent Reinforcement Learning

### v1

@@ Abstract
This paper studies networked multi-agent reinforcement learning (NMARL) with interdependent rewards and coupled policies. In this setting, each agent’s reward depends on its own state-action pair as well as those of its direct neighbors, and each agent’s policy is parameterized by its local parameters together with those of its $\kappa_{p}$-hop neighbors, with $\kappa_{p}\geq 1$ denoting the coupled radius. The objective of the agents is to collaboratively optimize their policies to maximize the discounted average cumulative reward. To address the challenge of interdependent policies in collaborative optimization, we introduce a novel concept termed the “neighbors’ averaged $Q$-function” and derive a new expression for the coupled policy gradient. Based on these theoretical foundations, we develop a distributed scalable coupled policy (DSCP) algorithm, where each agent relies only on the state-action pairs of its $\kappa_{p}$-hop neighbors and the rewards its their $(\kappa_{p}+1)$-hop neighbors. Specially, in the DSCP algorithm, we employ a geometric 2-horizon sampling method that does not require storing a full $Q$-table to obtain an unbiased estimate of the coupled policy gradient. Moreover, each agent interacts exclusively with its direct neighbors to obtain accurate policy parameters, while maintaining local estimates of other agents’ parameters to execute its local policy and collect samples for optimization. These estimates and policy parameters are updated via a push-sum protocol, enabling distributed coordination of policy updates across the network. We prove that the joint policy produced by the proposed algorithm converges to a first-order stationary point of the objective function. Finally, the effectiveness of DSCP algorithm is demonstrated through simulations in a robot path planning environment, showing clear improvement over state-of-the-art methods.
@@ V-B Main results
Suppose Assumptions 1-5 hold. In Algorithm 1, for all $i,j\in\mathcal{N}$, the local policy parameter estimate $\hat{\theta}^{i}_{j,t}$ of agent $i$ with respect to agent $j$ converges to agent $j$’s true policy parameter $\theta_{j,t}$ with the rate $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{O}(\frac{M_{\kappa_{p}}M_{\kappa_{p}+1}\log{t}}{t})$}, i.e.,
@@
V-C Extend results for reward function with larger coupled radius
Our proposed framework can also extend to settings where each agent’s reward depends on a larger neighborhood, under the following assumption.
@@ V-C Extend results for reward function with larger coupled radius
Under Assumption 6, the main results in Section V-B remain valid after adjusting the locality radii that appear in several key expressions. Specifically, the radius $(\kappa_{p}+1)$ in (1), (IV-A), and (13) is replaced by $(\kappa_{p}+\kappa_{r})$, and the radius $(\kappa_{p}+2)$ in (IV-A), (2), and (12) is replaced by $(\kappa_{p}+2\kappa_{r})$. Furthermore, Algorithm 1 requires redevelopment under the following assumption, which is a slightly modified version of Assumption 2.
@@
Suppose Assumptions 1 and 4-7 hold. In the extend Algorithm 1 for $\{r_{i}(s_{\mathcal{N}^{\kappa_{r}}_{i}},a_{\mathcal{N}^{\kappa_{r}}_{i}})\}_{i\in\mathcal{N}}$, the local policy parameter estimate $\hat{\theta}^{i}_{j,t}$ of agent $i$ with respect to agent $j$ converges to agent $j$’s true policy parameter $\theta_{j,t}$ with a rate $\mathcal{O}(\frac{M_{\kappa_{p}}M_{\kappa_{p}+\kappa_{r}}\log{t}}{t})$. Moreover, there exists a positive constant $J_{\mathrm{sup}}(M_{\kappa_{p}},M_{\kappa_{p}+\kappa_{r}})>0$ that is employed to substitute $J_{\mathrm{sup}}(M_{\kappa_{p}},M_{\kappa_{p}+1})$ in Theorem 4 to guarantee its validity.
@@ VI-A The description of the path planning problem
Figure 1: Path structure and underlying network.

### v2

@@ Abstract
This paper studies networked multi-agent reinforcement learning (NMARL) with interdependent rewards and coupled policies. In this setting, each agent’s reward depends on its own state-action pair as well as those of its direct neighbors, and each agent’s policy is parameterized by its local parameters together with those of its $\kappa_{p}$-hop neighbors, with $\kappa_{p}\geq 1$ denoting the coupled radius. The objective of the agents is to collaboratively optimize their policies to maximize the discounted average cumulative reward. To address the challenge of interdependent policies in collaborative optimization, we introduce a novel concept termed the “neighbors’ averaged $Q$-function” and derive a new expression for the coupled policy gradient. Based on these theoretical foundations, we develop a distributed scalable coupled policy (DSCP) algorithm, where each agent relies only on the state-action pairs of its $\kappa_{p}$-hop neighbors and the rewards of its $(\kappa_{p}+1)$-hop neighbors. Specially, in the DSCP algorithm, we employ a geometric 2-horizon sampling method that does not require storing a full $Q$-table to obtain an unbiased estimate of the coupled policy gradient. Moreover, each agent interacts exclusively with its direct neighbors to obtain accurate policy parameters, while maintaining local estimates of other agents’ parameters to execute its local policy and collect samples for optimization. These estimates and policy parameters are updated via a push-sum protocol, enabling distributed coordination of policy updates across the network. We prove that the joint policy produced by the proposed algorithm converges to a first-order stationary point of the objective function. Finally, the effectiveness of DSCP algorithm is demonstrated through simulations in a robot path planning environment, showing clear improvement over state-of-the-art methods.
@@ V-B Main results
Suppose Assumptions 1-5 hold. In Algorithm 1, for all $i,j\in\mathcal{N}$, the local policy parameter estimate $\hat{\theta}^{i}_{j,t}$ of agent $i$ with respect to agent $j$ converges deterministically to agent $j$’s true policy parameter $\theta_{j,t}$ with the rate $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{O}(\frac{M_{\kappa_{p}}M_{\kappa_{p}+1}\log{t}}{t})$}, i.e.,
@@
V-C Extended results for reward function with larger coupled radius
The proposed framework can be extended to settings where rewards depend on larger neighborhood scopes, as stated in the following assumption.
@@ V-C Extended results for reward function with larger coupled radius
It should be noted that the restriction $\kappa_{p}\geq\kappa_{r}$ in Assumption 6 is employed to guarantee that the coupled policy can comprehensively capture the inter-agent reward dependencies.
Under Assumption 6, the main results in Section V-B remain valid after appropriately adjusting the locality radii: $(\kappa_{p}+1)\to(\kappa_{p}+\kappa_{r})$ in (1), (IV-A), (13), and $(\kappa_{p}+2)\to(\kappa_{p}+2\kappa_{r})$ in (IV-A), (2), (12). Moreover, Algorithm 1 can then be similarly redeveloped under following assumption.
@@
Suppose Assumptions 1 and 4-7 hold. In the extended Algorithm 1 for $\{r_{i}(s_{\mathcal{N}^{\kappa_{r}}_{i}},a_{\mathcal{N}^{\kappa_{r}}_{i}})\}_{i\in\mathcal{N}}$, the local policy parameter estimate $\hat{\theta}^{i}_{j,t}$ of agent $i$ with respect to agent $j$ converges deterministically to agent $j$’s true policy parameter $\theta_{j,t}$ with a rate $\mathcal{O}(\frac{M_{\kappa_{p}}M_{\kappa_{p}+\kappa_{r}}\log{t}}{t})$. Moreover, there exists a positive constant $J_{\mathrm{sup}}(M_{\kappa_{p}},M_{\kappa_{p}+\kappa_{r}})>0$ that is employed to substitute $J_{\mathrm{sup}}(M_{\kappa_{p}},M_{\kappa_{p}+1})$ in Theorem 4 to guarantee its validity.
@@ VI-A The description of the path planning problem
Figure 1: The left panel depicts the path structure with 13 locations, while the right panel shows the underlying network among 10 agents.